# Be Your Engaged Partners

### v1

@@ Abstract
Large vision-language models (LVLMs), while proficient in following instructions and responding to diverse questions, invariably generate detailed responses even when questions are ambiguous or unanswerable, leading to hallucinations and bias issues. Thus, it is essential for LVLMs to proactively engage with humans to ask for clarifications or additional information for better responses. In this study, we aim to shift LVLMs from passive answer providers to proactive engaged partners. We begin by establishing a three-tiered hierarchy for questions of invalid, ambiguous, and personalizable nature to measure the proactive engagement capabilities of LVLMs. Utilizing this hierarchy, we create [Uncaptioned image]PIE (ProactIve Engagement Evaluation) through GPT-4o and human annotators, consisting of 853 questions across six distinct, fine-grained question types that are verified by human annotators and accompanied with well-defined metrics. Our evaluations on PIE indicate poor performance of existing LVLMs, with the best-performing open-weights model only achieving an Aggregate Align Rate (AAR) of 0.28. In response, we introduce [Uncaptioned image]MACAROON, self-iMaginAtion for ContrAstive pReference OptimizatiON, which instructs LVLMs to autonomously generate contrastive response pairs for unlabeled questions given the task description and human-crafted criteria. Then, the self-imagined data is formatted for conditional reinforcement learning. Experimental results show MACAROON effectively improves LVLMs’ capabilities to be proactively engaged (0.84 AAR) while maintaining comparable performance on general tasks [^1] The code will be made public at https://github.com/ShujinWu-0814/MACAROON..
@@
  Shujin Wu1,2∗    Yi R. Fung1    Sha Li1    Yixin Wan3     Kai-Wei Chang3    Heng Ji1 1University of Illinois Urbana-Champaign 2University of Southern California         3University of California, Los Angeles {shujinwu}@usc.edu         {yifung2, hengji}@illinois.edu
@@ 3.1 Self-Imagination
Constructing a preference dataset via human annotations is both resource-intensive and difficult to expand (Dai et al., 2024). Previous work also relies on proprietary LVLMs like GPT-4o for generate the golden responses, assuming that more advanced models would be available (Li et al., 2023c; Liu et al., 2023). In our study, we extend the “Constitutional AI” concept (Bai et al., 2022), and introduce a framework called “self-imagination”, which enables LVLMs to independently enhance their capabilities using human-defined criteria and unlabeled questions, which can be generated by a specific model or collected from the web demo.
@@ 6 Related Work
Multimodal pretraining has significantly advanced the landscape of vision-language tasks, with the emergence of unified pre-training frameworks designed to handle a diverse set of cross-modal and unimodal tasks Cho et al. (2021); Alayrac et al. (2022); Lu et al. (2022a). More recently, there has been increasing interest in visual instruction tuning Liu et al. (2023); Dai et al. (2024) as a pivotal methodology in the development of general-purpose LVLMs, enhancing their emergent in-context learning vision-language reasoning capabilities from zero-shot textual instruction and few-shot demonstration. Building upon this paradigm, later work further explores visual instruction tuning variations that incorporate better text reading localization Bai et al. (2023), OCR reading capability Liu et al. (2024), object attribute relations Zhao et al. (2022), open world knowledge Liu et al. (2024), and efficiency considerations Hu et al. (2024). Nevertheless, LVLMs are prone to unexpected behaviors that may not align well with human intents Qiu et al. (2024). Towards this end, there has been work such as Llava-Guard for ensuring the safety compliance of visual content against toxicity or violent-prone threats Helff et al. . However, directly mitigating multimodal hallucination is largely an unexplored research area that our work proposes to address. We further discuss related work on efforts to reduce hallucination and understand intents for LLMs in Appendix C due to space limits.

### v2

@@ Abstract
Large vision-language models (LVLMs), while proficient in following instructions and responding to diverse questions, invariably generate detailed responses even when questions are ambiguous or unanswerable, leading to hallucinations and bias issues. Thus, it is essential for LVLMs to proactively engage with humans to ask for clarifications or additional information for better responses. In this study, we aim to shift LVLMs from passive answer providers to proactive engaged partners. We begin by establishing a three-tiered hierarchy for questions of invalid, ambiguous, and personalizable nature to measure the proactive engagement capabilities of LVLMs. Utilizing this hierarchy, we create [Uncaptioned image]PIE (ProactIve Engagement Evaluation) through GPT-4o and human annotators, consisting of 853 questions across six distinct, fine-grained question types that are verified by human annotators and accompanied with well-defined metrics. Our evaluations on PIE indicate poor performance of existing LVLMs, with the best-performing open-weights model only achieving an Aggregate Align Rate (AAR) of 0.28. In response, we introduce [Uncaptioned image]MACAROON, self-iMaginAtion for ContrAstive pReference OptimizatiON, which instructs LVLMs to autonomously generate contrastive response pairs for unlabeled questions given the task description and human-crafted criteria. Then, the self-imagined data is formatted for conditional reinforcement learning. Experimental results show MACAROON effectively improves LVLMs’ capabilities to be proactively engaged (0.84 AAR) while maintaining comparable performance on general tasks [^1] The code is made public at https://github.com/ShujinWu-0814/MACAROON..
@@
  Shujin Wu1,2∗    May Fung1    Sha Li1    Yixin Wan3     Kai-Wei Chang3    Heng Ji1 1University of Illinois Urbana-Champaign 2University of Southern California         3University of California, Los Angeles {shujinwu}@usc.edu         {yifung2, hengji}@illinois.edu
@@ 3.1 Self-Imagination
Constructing a preference dataset via human annotations is both resource-intensive and difficult to expand (Dai et al., 2024). Previous work also relies on proprietary LVLMs like GPT-4o to generate the golden responses, assuming that more advanced models would be available (Li et al., 2023c; Liu et al., 2023). In our study, we extend the “Constitutional AI” concept (Bai et al., 2022), and introduce a framework called “self-imagination”, which enables LVLMs to independently enhance their capabilities using human-defined criteria and unlabeled questions, which can be generated by a specific model or collected from the web demo.
@@ 6 Related Work
Multimodal pretraining has significantly advanced the landscape of vision-language tasks, with the emergence of unified pre-training frameworks designed to handle a diverse set of cross-modal and unimodal tasks Cho et al. (2021); Alayrac et al. (2022); Lu et al. (2022a). More recently, there has been increasing interest in visual instruction tuning Liu et al. (2023); Dai et al. (2024) as a pivotal methodology in the development of general-purpose LVLMs, enhancing their emergent in-context learning vision-language reasoning capabilities from zero-shot textual instruction and few-shot demonstration. Building upon this paradigm, later work further explores visual instruction tuning variations that incorporate better text reading localization Bai et al. (2023), OCR reading capability Liu et al. (2024), object attribute relations Zhao et al. (2022), open world knowledge Liu et al. (2024), and efficiency considerations Hu et al. (2024). Nevertheless, LVLMs are prone to unexpected behaviors that may not align well with human intents Qiu et al. (2024); Wang et al. (2024). Towards this end, there has been work such as Llava-Guard for ensuring the safety compliance of visual content against toxicity or violent-prone threats Helff et al. . However, directly mitigating multimodal hallucination is largely an unexplored research area that our work proposes to address. We further discuss related work on efforts to reduce hallucination and understand intents for LLMs in Appendix C due to space limits.
@@ 7 Conclusions
Acknowledgement
This research is based upon work supported DARPA ITM Program No. FA8650-23-C-7316 and the AI Research Institutes program by National Science Foundation and the Institute of Education Sciences, U.S. Department of Education through Award # 2229873 - AI Institute for Transforming Education for Children with Speech and Language Processing Challenges. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.